Session: Mar 2008 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Fri Feb 29 13:33:21 2008) Success :
No known problems.

Urumqi (Hua Zhang at Fri Feb 29 13:37:00 2008) Success :
No known problems.

Onsala (Michael Lindqvist at Fri Feb 29 13:58:16 2008) Success :
Formatter to fs time difference. Fixed after first scan.

Medicina (Morsiani at Fri Feb 29 14:05:48 2008) Success :

Effelsberg (Uwe Bach at Mon Mar 3 12:15:43 2008) Minor failures : RECO
Lost 20 minutes (scans 52-57) due to connection problems during a disk-pack change.

Jodrell_Bank (Alastair Gunn at Mon Mar 3 15:46:35 2008) Success :
The Lovell Telescope was used (Jb1). No known problems.

Cambridge (Alastair Gunn at Mon Mar 3 15:47:41 2008) Minor failures : RECO
Problem with data throttling. Tried diskpack swaps after bank change but did not resolve the problem. Swapped to spare formatter AFTER this experiment.

Noto (P. Cassaro at Tue Mar 4 17:09:23 2008) Minor failures : TSYS
The LSB of the bbc #8 was in overflow.

Westerbork (Geert Kuper at Mon Mar 17 15:42:42 2008) Minor failures :
Wb used in phased array mode. Wb used in phased array mode. One of the disk packs has the wrong VSN; in the log file it says OSOD-057 but the external label is JIVE-020